OpenStudy (anonymous):

find the x and y intercepts of the line that is tangent to the curve y=x^3 at the point (-3,-27)

OpenStudy (anonymous):

differentiate Y'=3x^2 slope at that point is 3(-3)^2=27 equation of tgt is Y-y=m(X-x) Y-27X=54 to get intercepts hide Y xintercept is-2.thatis the point (-2,0) hide X ,Yintercept is54.that is the point (0,54)